OpenStudy (hanner_b_nanner):

Fan&Medal... Can someone check my answer? Determine if triangle ABC with coordinates A (−1, 1), B (1, −1), and C (3, 3) is an isosceles triangle. Use evidence to support your claim. If it is not an isosceles triangle, what changes can be made to make it isosceles? Be specific.

OpenStudy (hanner_b_nanner):

You can use the distance formula to find out the lengths of all three sides. If two of the sides are the same, the triangle will be an isosceles. AC d=(x2-x1)^2 +(y2-y1)^2 d=(3+1)^2+(3-1)^2 d=(4)^2+(2)^2 d=(16)+(4)=sqrt20 ----------------------- AB d=(1+1)^2 + (-1-1)^2 d=(2)^2+(-2)^2 d=(4)+(4)=sqrt8 ----------------------- BC d=(3-1)^2+(3+1)^2 d=(2)^2+(4)^2 d=(4)+(16)=sqrt20 So yes, this triangle is an isosceles triangle because two of the sides equal 20.

OpenStudy (princeharryyy):

thumbs up !!!! @hanner_B_nanner

OpenStudy (princeharryyy):

u there ???? @hanner_B_nanner

OpenStudy (hanner_b_nanner):

yesss!! Thank you!!!! :D

OpenStudy (princeharryyy):

u r welcome .... @hanner_B_nanner